what is the molar solubility of Zn^2+ in a solution that is buffered at a pH of 3.00?

The Ksp for Zn(OH)2 is 3x10^-17.
Answer #1

pOH = 14 - pH

= 14 - 3.00 = 11.0

OH- = 10^-pOH

= 10^ -11

Zn(OH)2(s) = Zn2+(aq) + 2OH-(aq)

S 10^-11

Ksp = (Zn2+)*(OH-)^2

= 3*10^-17 = S * ( 10^-11)^2

S = 3* 10^5 M

If you have any questions please comment

If you satisfied with the solution please rate it thanks